# Volatility Parametrizations with Random Coefficients: Analytic Flexibility for Implied Volatility Surfaces

### v2

@@ 4.1 Randomized Spot Prices
Suppose that $\hat{\sigma}(T,K,\overline{p})$ is a parametrization with parameters $\overline{p}$ and let $p_{S_{T};\overline{p}}$ be the risk-neutral PDF given by the parametrization for an asset $S$ at the fixed time $T$. Since the risk-neutral probability density is a PDF of a random variable $S_{T}$, a necessary condition for it to be arbitrage-free is that $S_{T}$ is centered at its forward $S_{0}{\rm e}^{r(T-t_{0})}$. This can be shown as if
@@
then, an integration of Equation 4.1 multiplied with $x$ shows that the expectation of $S_{T}$ is $\mathbb{E}[S_{T}]=S_{0}{\rm e}^{r(T-t_{0})}$. If we assume $\vartheta$ to be an absolutely continuous random variable with mean $\mathbb{E}[\vartheta]=S_{0}$ and assume it is almost surely positive, we can consider a randomization of the spot price by replacing $S_{0}$ with the random variable $\vartheta$. As it will be presented later, the fact that it is centered at $S_{0}$ is necessary to ensure that the randomization is arbitrage-free. Denoting $f_{\vartheta}(x)$ the PDF of $\vartheta$, we find equivalently to Equation 2.7:
$\mathbb{E}[V_{c/p}(\vartheta,T,K;\overline{p})]=\int_{\mathbb{R}}BS_{c/p}(t_{0},\theta,T,K;\hat{\sigma}(T,K;\overline{p}))f_{\vartheta}(\theta){\rm d}\theta,$ (4.2)
where the value of the call/put options is now denoted by $V_{c/p}(S_{0},T,K;\overline{p})$ with the extra parameter $S_{0}$. The randomized price surface Equation 4.2 is an average of Black-Scholes prices, but the varying parameter is the spot price, as opposed to the parameters $\overline{p}$ as before. Ensuring that the spot randomization is given in terms of a set of parameters, we define an extended parameter vector $\overline{p}^{*}=(p_{1},p_{2},\dots,p_{m},q_{1},q_{2},\dots)\in\mathcal{D}^{*}$, which contains all previous parameters plus the parameters $(q_{1},q_{2}\dots)$ specifying the distribution of $\vartheta$. This defines the pricing surface of the spot randomization, and we can define the randomized volatility surface $\hat{\sigma}_{s}(T,K;\overline{p}^{*})$ with the new parameter set as the Black-Scholes inverse of Equation 4.2.
@@
Let $\vartheta$ be centered at $S_{0}$, absolutely continuous and almost surely positive. Then, the randomized spot volatility surface $\hat{\sigma}_{s}(T,K;\overline{p}^{*})$, given by Equation 4.2, is arbitrage-free.
@@
We will show that we can write $\mathbb{E}[V_{c/p}(\vartheta,T,K;\overline{p})]$ as an integral of a proper probability density function centered at $S_{0}{\rm e}^{r(T-t_{0})}$, which is sufficient to ensure the absence of arbitrage [4]. Suppose that $p_{S_{T};\theta}$ is the PDF according to Equation 4.1, where $S_{0}$ is replaced with $\theta$. In this case, we obtain
$\mathbb{E}[V_{c/p}(\vartheta,T,K;\overline{p})]$ $=\int_{\mathbb{R}}BS_{c/p}(t_{0},\theta,T,K;\hat{\sigma}_{s}(T,K;\overline{p}))f_{\vartheta}(\theta){\rm d}\theta$
$=\int_{\mathbb{R}}\int_{K}^{\infty}(x-K)p_{S_{T};\theta}(x)f_{\vartheta}(\theta){\rm d}x{\rm d}\theta$
$=\int_{K}^{\infty}(x-K)\int_{\mathbb{R}}p_{S_{T};\theta}(x)f_{\vartheta}(\theta){\rm d}\theta{\rm d}x$
@@
$\int_{0}^{\infty}x\mathbb{E}\left[p_{S_{T};\vartheta}(x)\right]{\rm d}x$ $=\int_{0}^{\infty}x\int_{\mathbb{R}}p_{S_{T};\theta}(x)f_{\vartheta}(\theta){\rm d}\theta{\rm d}x$
$=\int_{\mathbb{R}}\int_{0}^{\infty}xp_{S_{T};\theta}(x){\rm d}xf_{\vartheta}(\theta){\rm d}\theta$
$=\int_{\mathbb{R}}\theta{\rm e}^{r(T-t_{0})}f_{\vartheta}(\theta){\rm d}\theta$
@@
It is apparent that as long as $\vartheta$ is centered at $S_{0}$, the randomization does not induce any butterfly arbitrage, as the forward price, whose PDF is given by $\mathbb{E}\left[p_{S_{T};\vartheta}(x)\right]$, is centered at $S_{0}{\rm e}^{r(T-t_{0})}$. We will now continue the randomization process by discretizing the integral using the quadrature method. We define the discrete randomized pricing surface of the spot price as
@@
Suppose we obtain the risk-neutral probability density function $f_{S_{T}}(\cdot)$ of the asset price at a fixed expiry $T$. Consider a spot randomization of the parametric surface $\hat{\sigma}(T,K,\sigma)=\sigma$ with $\sigma=0$ and random spot price given by the PDF of $f_{S_{T}}(\cdot)$. If $\hat{\sigma}(T,K,\sigma)=0$, the call option price $V_{c}(S_{0},T,K,\sigma)$ is given by $(S_{0}-K)^{+}$, which means that
@@
The right-hand side is exactly the call option price function given the risk-neutral probability density function $f_{S_{T}}(\cdot)$. The randomization therefore perfectly prices any option with expiry $T$.
@@ 4.2 Illustrative Example: Spot Randomization of Flat Surface
We will again first consider the simplest randomization, which is the flat volatility surface $\hat{\sigma}(T,K;\sigma)=\sigma$. Instead of replacing $\sigma$ with a random variable, we randomize the spot price $S_{0}$ using a lognormal distributed random variable $\vartheta$, such that $\log(\vartheta)\sim\mathcal{N}(\log(S_{0})-\frac{\nu^{2}}{2},\nu^{2})$ given the parameter $\nu$. The particular choice for the average ensures that
@@
and therefore, that the randomization is arbitrage-free. We compute the prices, implied volatilities, and risk-neutral probability densities for various values of $\nu$. The results are shown in Figure 7. The third graph shows that the risk-neutral probability density functions exhibits a bimodal shape, while the implied volatility is strongly concave. Next to the exact implied volatilities, we also plot 2nd to 4th-order approximations for the three shapes and highlight the convergence of the approximation to the exact implied volatility curve, which is derived using a root-finding scheme.
Figure 7: PDF and Implied Volatility of a randomized spot flat volatility surface. The implied volatility is derived from a root finder in the first few orders of approximations. The graph shows the convergence of the expansion to the reference graph.
@@ 4.3 Fitting an Earnings Announcement Volatility Surface
To demonstrate the capabilities of the randomization of the spot price, particularly in regard to earnings announcements, we consider an example of real market quotes. The company Amazon.com Inc. announced its earnings release for 2018 Q1 after market close on Apr 26th, 2018, which induced uncertainty in the market on the day of Apr 26th. We obtained the market quotes for options traded on the stock on the morning of the 26th (10:30 AM), which expire the next day on Apr 27th. Table 5 shows the summary statistics of the quotes which we obtain from the Cboe Data Shop [5].
@@
Using the quotes, we fit a randomized spot SABR parametrization on the options quotes, where the spot price $S_{0}$ is again randomized with a log-normally distributed random variable $\vartheta$, such that $\log\vartheta\sim\mathcal{N}(\log(S_{0})-\frac{\nu^{2}}{2},\nu^{2})$. For the discretization, we choose $N_{q}=2$ and run an optimization algorithm to obtain the optimal parameters for $\overline{p}^{*}=(\beta,\alpha,\rho,\gamma,\nu)$. As a benchmark, we also fit an SVI-type parametrization, which is not randomized, to the data. Figure 8 shows the fit of the implied volatility parametrizations, the market quotes, and the risk-neutral density from the randomized parametrization. The randomized spot parametrization is able to reproduce the shape of the implied volatility of the quotes. Furthermore, we see from the risk-neutral density this volatility shape stems from a bimodal probability density. The red line in the plot indicates the strike at the mean of the distribution (which is the forward-ATM). The SVI parametrization, on the other hand, fails to reproduce the implied volatility shape entirely, fitting to an essentially straight line.
Figure 8: Risk-neutral PDF and implied volatilities of randomized parametrization.
@@ 5 Conclusion
This paper introduces a novel generic method to enhance the flexibility of volatility surface parametrizations through parameter randomization. We formulated new parametric surfaces from existing ones by replacing one of the parameters with a random variable. The method first defines a distribution for a set of parameters and then formulates the randomization as the expectation of the European option price under the distribution. This induces a mixture-type behavior in the pricing surface of the options and through Breeden-Litzenberger [3] also in the risk-neutral probability density functions. The mixture increases the flexibility of the price and, thus, provides a better ability to fit the market quotes when traditional parametrizations fail. Lastly, we derive an expansion of the implied volatility surface as a function of the input parameters, which can be computed to an arbitrary degree of accuracy. We presented two examples of randomized surfaces and showed that the randomized SABR parametrization is able to fit the data better than a classical SABR parametrization or an SVI-type parametrization.
In the second part of the paper, we utilized the randomization technique to formulate a randomized spot volatility surface. With this method, the spot price $S_{0}$ of the asset is randomized using a random variable centered at $S_{0}$. This type of randomization is particularly effective in fitting volatility shapes that imply risk-neutral distribution functions of multi-modal form, which often occur shortly prior to earnings announcements of equities. Again, we derived an expansion of the implied volatility based on the same techniques as before and showed that the parametrization fits well with the data for near-maturity options during earnings announcements, a task that is impossible for classical diffusion-based parametrizations.
@@ Appendix A Quadrature Pairs $\{\lambda_{n},x_{n}\}_{n\leq N_{q}}$ for Expectations
Here, we present the derivation of the quadrature points to calculate the expectation of the randomized pricing surface. The section is based on [13] and Golub and Welsch [11], which establish the algorithm for the computation of the quadrature points for random variables based on the moments.
The Gaussian quadrature points enable an efficient approximation of the expectation $\mathbb{E}[g(X)]$ given an arbitrary function $g(\cdot)$ and an absolutely continuous random variable $X$, such that $f_{X}(x)$ is its probability density function. Since the expectation is an integral over the domain $\mathcal{D}$ of $X$, the integral can be written as:
$\mathbb{E}[g(X)]=\int_{\mathcal{D}}g(x)f_{X}(x){\rm d}x\approx\sum_{n=1}^{N_{q}}\lambda_{n}g(x_{n}),$ (A.1)
where $\{\lambda_{n},x_{n}\}_{n\leq N_{q}}$ points are chosen in an optimal way, which we will establish below. The algorithm is general for any type of expectation and requires only an efficient computation of the moments $\mu_{i}=\mathbb{E}[X^{i}]$ of $X$. The foundation of the algorithm is to establish a sequence of polynomials $p_{0}(x),p_{1}(x),\dots$, which are orthonormal with respect to $X$, i.e., such that
$\mathbb{E}[p_{i}(X)p_{j}(X)]=\begin{cases}1\quad&\text{ if }\quad i=j,\\ 0\quad&\text{ otherwise. }\end{cases}$ (A.2)
@@
and additionally the quadrature weights $\lambda_{n}=(v^{n}_{1})^{2}$ by the square of the first row of the $n$-th eigenvalue $v_{n}$ (see [11] for an extensive discussion). An implementation of the algorithm in Python and MATLAB can be found on Github [12].
Appendix B Higher-Order (3rd,4th) Terms Calculation for Spot Randomization

### v3

@@ 4.1 Randomized Spot Parameter
Let $S_{0}$ be the current spot price of a financial asset $S$ at time $t_{0}$. Suppose that $\hat{\sigma}(T,K,\overline{p})$ is a parametrization with parameters $\overline{p}$ and let $p_{S_{T};\overline{p}}$ be the probability density function under a suitable risk-neutral measure $\mathbb{Q}$, given by the parametrization for $S$ at the fixed time $T$. Since the risk-neutral probability density is a PDF of a random variable $S_{T}$ on a risk-neutral probability space, a necessary condition for it to be arbitrage-free is that $S_{T}$ is centered at its forward $S_{0}{\rm e}^{r(T-t_{0})}$. This can be shown as if
@@
then, an integration of Equation 4.1 multiplied with $x$ shows that the expectation of $S_{T}$ is $\mathbb{E}^{\mathbb{Q}}[S_{T}]=S_{0}{\rm e}^{r(T-t_{0})}$. If we assume $\vartheta$ to be a random variable with law $\mu$ and mean $\mathbb{E}[\vartheta]=S_{0}$ and assume it is almost surely positive, we can consider a randomization of the spot entry $s$ of the Black-Scholes formula $BS_{c/p}(t_{0},s,T,K;\hat{\sigma}(T,K))$, by replacing $s$ with the random variable $\vartheta$. As it will be presented later, the fact that it is centered at $S_{0}$ is necessary to ensure that the randomization is arbitrage-free. We find equivalently to Equation 2.7:
$\mathbb{E}[V_{c/p}(\vartheta,T,K;\overline{p})]=\int_{\mathbb{R}}BS_{c/p}(t_{0},\theta,T,K;\hat{\sigma}(T,K;\overline{p})){\rm d}\mu(\theta),$ (4.2)
where the value of the call/put options is now denoted by $V_{c/p}(S_{0},T,K;\overline{p})$ with the extra parameter $S_{0}$. The randomized price surface Equation 4.2 is an average of Black-Scholes prices, where the varying entry is the spot price entry, as opposed to the parameters $\overline{p}$ as before. Ensuring that the spot randomization is given in terms of a set of parameters, we define an extended parameter vector $\overline{p}^{*}=(p_{1},p_{2},\dots,p_{m},q_{1},q_{2},\dots)\in\mathcal{D}^{*}$, which contains all previous parameters plus the parameters $(q_{1},q_{2}\dots)$ specifying the distribution of $\vartheta$. This defines the pricing surface of the spot randomization, and we can define the randomized volatility surface $\hat{\sigma}_{s}(T,K;\overline{p}^{*})$ with the new parameter set as the Black-Scholes inverse of Equation 4.2.
@@
Let $\vartheta$ be centered at $S_{0}$ and almost surely positive. Then, the randomized spot volatility surface $\hat{\sigma}_{s}(T,K;\overline{p}^{*})$, given by Equation 4.2, is arbitrage-free.
@@
We will show that we can write $\mathbb{E}[V_{c/p}(\vartheta,T,K;\overline{p})]$ as an integral of a proper probability density function centered at $S_{0}{\rm e}^{r(T-t_{0})}$, which is sufficient to ensure the absence of arbitrage [brigo2002lognormal]. Suppose that $p_{S_{T};\theta}$ is the PDF according to Equation 4.1, where $S_{0}$ is replaced with $\theta$. In this case, we obtain
$\mathbb{E}[V_{c/p}(\vartheta,T,K;\overline{p})]$ $=\int_{\mathbb{R}}BS_{c/p}(t_{0},\theta,T,K;\hat{\sigma}_{s}(T,K;\overline{p})){\rm d}\mu(\theta)$
$=\int_{\mathbb{R}}\int_{K}^{\infty}(x-K)p_{S_{T};\theta}(x){\rm d}x{\rm d}\mu(\theta)$
$=\int_{K}^{\infty}(x-K)\int_{\mathbb{R}}p_{S_{T};\theta}(x){\rm d}\mu(\theta){\rm d}x$
@@
$\int_{0}^{\infty}x\mathbb{E}\left[p_{S_{T};\vartheta}(x)\right]{\rm d}x$ $=\int_{0}^{\infty}x\int_{\mathbb{R}}p_{S_{T};\theta}(x){\rm d}\mu(\theta){\rm d}x$
$=\int_{\mathbb{R}}\int_{0}^{\infty}xp_{S_{T};\theta}(x){\rm d}x{\rm d}\mu(\theta)$
$=\int_{\mathbb{R}}\theta{\rm e}^{r(T-t_{0})}{\rm d}\mu(\theta)$
@@
This shows that as long as $\vartheta$ is centered at $S_{0}$, the randomization is free of arbitrage, as the forward price, whose PDF is given by $x\mapsto\mathbb{E}\left[p_{S_{T};\vartheta}(x)\right]$, is centered at $S_{0}{\rm e}^{r(T-t_{0})}$. Note that the spot randomization does not model any uncertainty of the spot price $S_{0}$, which remains a fixed, deterministic input to the model. The randomization is merely defined as randomizing the spot parameter $s$ of the Black-Scholes formula $BS_{c/p}(t_{0},s,T,K;\hat{\sigma}(T,K))$. Intuitively, the randomization introduces uncertainty about potential future trends in the underlying asset’s price evolution. Since the Black-Scholes formula implies a diffusion model centered at the forward price (which depends on $r$ and the spot parameter), the risk-neutral PDF of the Black-Scholes model assumes the underlying price at expiry to be centered on the forward price, with symmetric uncertainty around this expectation. This implies the model assigns the highest probability to minimal price movements, such that the mode and mean of the risk-neutral probability density coincide. Such models struggle to accommodate scenarios where large price movements are anticipated, but the direction (up or down) remains uncertain. By randomizing the spot parameter input to the Black-Scholes formula, we create a mixture of multiple Black-Scholes dynamics: the resulting risk-neutral density can become multimodal, with peaks representing distinct market views on possible future mean prices. Crucially, this randomization does not affect the $S_{0}$, but instead reflects uncertainty about where the price process will center itself in the future. The framework preserves arbitrage-free pricing, as the expected value of the future spot price still matches the forward price.
We now continue the randomization process by discretizing the integral using the quadrature method. We define the discrete randomized pricing surface of the spot parameter as
@@
Suppose we obtain the risk-neutral probability density function $f_{S_{T}}(\cdot)$ of the asset price at a fixed expiry $T$. Consider a spot randomization of the parametric surface $\hat{\sigma}(T,K,\sigma)=\sigma$ with $\sigma=0$ and random spot parameter given by the PDF of $f_{S_{T}}(\cdot)$. If $\hat{\sigma}(T,K,\sigma)=0$, the call option price $V_{c}(S_{0},T,K,\sigma)$ is given by $(S_{0}-K)^{+}$, which means that
@@
The right-hand side is exactly the call option price function given the risk-neutral probability density function $f_{S_{T}}(\cdot)$. The randomization perfectly prices any option with expiry $T$.
@@ 4.2 Illustrative Example: Spot Randomization of Flat Surface
We again first consider the simplest randomization, which is the flat volatility surface $\hat{\sigma}(T,K;\sigma)=\sigma$. Instead of replacing $\sigma$ with a random variable, we randomize the spot parameter using a lognormal distributed random variable $\vartheta$, such that $\log(\vartheta)\sim\mathcal{N}(\log(S_{0})-\frac{\nu^{2}}{2},\nu^{2})$ given the parameter $\nu$. The particular choice for the average ensures that
@@
and therefore, that the randomization is arbitrage-free. We compute the prices, implied volatilities, and risk-neutral probability densities for various values of $\nu$ with $N_{q}=2$. The other parameters are $S_{0}=3,T=1,\sigma=0.12,r=0\%$ and the results are shown in Figure 8. The third graph shows that the risk-neutral probability density functions exhibit a bimodal shape, while the implied volatility is strongly concave. Next to the exact implied volatilities, we also plot 2nd to 4th-order approximations for the three shapes and highlight the convergence of the approximation to the exact implied volatility curve, which is derived using a root-finding scheme.
Figure 8: PDF and Implied Volatility of a randomized spot flat volatility surface. The implied volatility is derived from a root finder in the first few orders of approximations. The graph shows the convergence of the expansion to the reference graph. The mixture parameters are $(\lambda_{1},\lambda_{2},\theta_{1},\theta_{2})=(0.57,0.43,2.74,3.35),(0.61,0.39,2.64,3.57)$ and $(0.64,0.36,2.55,3.82)$ from left to right.
@@ 4.3 Fitting an Earnings Announcement Volatility Surface
To demonstrate the capabilities of the randomization of the spot parameter, particularly in regard to earnings announcements, we consider an example of real market quotes. The company Amazon.com Inc. announced its earnings release for 2018 Q1 after market close on Apr 26th, 2018, which induced uncertainty in the market on the day of Apr 26th. We obtained the market quotes for options traded on the stock on the morning of the 26th (10:30 AM), which expire the next day on Apr 27th. Table 5 shows the summary statistics of the quotes which we obtain from the Cboe Data Shop [cboeSPX].
@@
Using the quotes, we fit a randomized spot SABR parametrization on the options quotes, where the spot parameter is randomized with a log-normally distributed random variable $\vartheta$, such that $\log\vartheta\sim\mathcal{N}(\log(S_{0})-\frac{\nu^{2}}{2},\nu^{2})$. For the discretization, we choose $N_{q}=2$ and run an optimization algorithm to obtain the optimal parameters for $\overline{p}^{*}=(\beta,\alpha,\rho,\gamma,\nu)$. As a benchmark, we also fit a SABR and SVI-type parametrization, which are not randomized, to the data. Furthermore, we also fit a multimodal lognormal mixture parametrization, which is designed to produce multimodal mixture densities. Figure 9 shows the fit of the implied volatility parametrizations, the market quotes, and the risk-neutral density from the randomized parametrization. The randomized spot parametrization is able to reproduce the shape of the implied volatility of the quotes. Furthermore, we see from the risk-neutral density this volatility shape stems from a bimodal probability density. The red line in the plot indicates the strike at the mean of the distribution (which is the forward-ATM). The SVI/SABR parametrizations, on the other hand, fail to reproduce the implied volatility shape entirely, fitting to an essentially straight line.
Figure 9: Risk-neutral PDF and implied volatilities of randomized parametrization. The fitted parameters for the randomized SABR are $(\beta,\alpha,\rho,\gamma,\nu)=(0.9,1.38,-0.03,10.99,0.46)$. This is equivalent to $\theta_{1},\theta_{2}=1432,1570$ with weights $\lambda_{1},\lambda_{2}=0.53,0.47$.
@@ 5 Conclusion
This paper introduces a novel generic method to enhance the flexibility of volatility surface parametrizations through parameter randomization. We formulated new parametric surfaces from existing ones by replacing one of the parameters with a random variable. The method first defines a distribution for a set of parameters and then formulates the randomization as the expectation of the European option price under the distribution. This induces a mixture-type behavior in the pricing surface of the options and through Breeden-Litzenberger [breeden1978prices] also in the risk-neutral probability density functions. The mixture increases the flexibility of the price and, thus, provides a better ability to fit the market quotes when traditional parametrizations fail. Lastly, we derive an expansion of the implied volatility surface as a function of the input parameters, which can be computed to an arbitrary degree of accuracy. We presented two examples of randomized surfaces and showed that the randomized SABR parametrization is able to fit the data better than a classical SABR parametrization or an SVI-type parametrization.
In the second part of the paper, we utilized the randomization technique to formulate a randomized spot volatility surface. With this method, the spot parameter of the Black-Scholes formula is randomized using a random variable centered at the current spot price $S_{0}$. This type of randomization is particularly effective in fitting volatility shapes that imply risk-neutral distribution functions of multi-modal form, which often occur shortly prior to earnings announcements of equities. Again, we derived an expansion of the implied volatility based on the same techniques as before and showed that the parametrization fits well with the data for near-maturity options during earnings announcements, a task that is impossible for classical diffusion-based parametrizations.
@@ Appendix A Quadrature Pairs $\{\lambda_{n},x_{n}\}_{n\leq N_{q}}$ for Expectations
Here, we present the derivation of the quadrature points to calculate the expectation of the randomized pricing surface. The section is based on [grzelak2019stochastic] and Golub and Welsch [golub1969calculation], which establish the algorithm for the computation of the quadrature points for random variables based on the moments.
The Gaussian quadrature points enable an efficient approximation of the expectation $\mathbb{E}[g(X)]$ given an arbitrary function $g(\cdot)$ and a random variable $X$ with law $\mu$. Since the expectation is an integral over the domain $\mathcal{D}$ of $X$, the integral can be written as:
$\mathbb{E}[g(X)]=\int_{\mathcal{D}}g(x){\rm d}\mu(x)\approx\sum_{n=1}^{N_{q}}\lambda_{n}g(x_{n}),$ (A.1)
where $\{\lambda_{n},x_{n}\}_{n\leq N_{q}}$ points are chosen in an optimal way, which we establish below. The algorithm is general for any type of expectation and requires only an efficient computation of the moments $\mu_{i}=\mathbb{E}[X^{i}]$ of $X$. The foundation of the algorithm is to establish a sequence of polynomials $p_{0}(x),p_{1}(x),\dots$, which are orthonormal with respect to $X$, i.e., such that
[…]
@@
and additionally the quadrature weights $\lambda_{n}=(v^{n}_{1})^{2}$ by the square of the first row of the $n$-th eigenvalue $v_{n}$ (see [golub1969calculation] for an extensive discussion). An implementation of the algorithm in Python and MATLAB can be found on GitHub [github].
Appendix B Supplementary Results for SPX Fitting
In this appendix, we provide additional information and results for the SPX fitting experiment.
Table 6 summarizes the data of the option chains and expiries available in the set. Included in the set were all options with a valid bid quote at the EOD snapshot for the main expiry dates. When both put and call quotes are available we select the most liquid quote per strike based on the open interest, which are the out-the-money quotes. The quotes were not cleaned for arbitrage before fitting.
Expiry Date Days to Maturity N Quotes Min Strike Max Strike
16-Aug-2024 16 364 3225 6140
20-Sep-2024 51 352 3150 7100
18-Oct-2024 79 339 3000 8000
15-Nov-2024 107 289 3000 8400
20-Dec-2024 142 233 2400 9000
17-Jan-2025 170 228 2000 9000
21-Feb-2025 205 147 1800 9000
21-Mar-2025 233 148 1400 9000
20-Jun-2025 324 160 1400 9000
19-Sep-2025 415 144 1400 8800
19-Dec-2025 506 144 1400 12000
18-Dec-2026 870 88 1400 12000
17-Dec-2027 1234 58 1400 12000
Table 6: Data summary implied volatilities quotes on Jul 31st, 2024. The strike range is chosen based on sufficiently liquid quotes available.
Figure 10 and Figure 11 show the fitting results for the randomized SABR vs the two benchmarks, and Table 7, Table 8 show the mean squared errors. In both cases, the randomized SABR performs unanimously the best, matching the market quotes closely.
Figure 10: Option Chain for Aug 30 2024
Figure 11: Option Chain for Sep 30 2024
Aug 30 2024 Mean Squared Errors (MSE)
Expiry Month Lognormal Mixture Regular SABR Randomized SABR
Sep 24 1.14E-03 1.68E-04 7.59E-05
Oct 24 1.05E-03 4.31E-05 2.54E-05
Nov 24 1.08E-03 4.62E-05 1.70E-05
Dec 24 1.05E-03 5.62E-05 6.90E-06
Jan 25 1.27E-03 7.16E-05 8.20E-06
Feb 25 1.22E-03 8.59E-05 5.20E-06
Mar 25 1.81E-03 6.89E-05 5.10E-06
Jun 25 1.43E-03 1.19E-04 1.90E-06
Sep 25 1.08E-03 1.67E-04 2.40E-06
Dec 25 8.33E-04 1.91E-04 6.00E-06
Dec 26 8.60E-04 3.24E-04 1.04E-05
Dec 27 7.20E-04 3.42E-04 1.01E-05
Dec 28 4.62E-04 5.01E-04 6.07E-05
Dec 29 2.31E-04 5.83E-04 7.00E-06
Table 7: Mean Square Error across entire dataset - Aug 30
Sep 30 2024 Mean Squared Errors (MSE)
Expiry Month Lognormal Mixture Regular SABR Randomized SABR
Oct 24 1.24E-03 2.97E-04 7.41E-05
Nov 24 1.78E-03 9.29E-05 5.65E-05
Dec 24 1.74E-03 7.08E-05 3.00E-05
Jan 25 1.45E-03 9.37E-05 8.00E-06
Feb 25 1.40E-03 9.36E-05 7.30E-06
Mar 25 1.47E-03 1.04E-04 5.10E-06
Jun 25 1.70E-03 1.43E-04 5.70E-06
Sep 25 1.18E-03 1.83E-04 1.84E-05
Dec 25 8.64E-04 2.10E-04 1.44E-05
Dec 26 1.01E-03 3.30E-04 1.90E-05
Dec 27 8.44E-04 4.01E-04 3.37E-05
Dec 28 5.25E-04 5.29E-04 9.87E-05
Dec 29 2.22E-04 6.17E-04 6.20E-06
Table 8: Mean Square Error across entire dataset - Sep 30
Appendix C Higher-Order (3rd,4th) Terms Calculation for Spot Randomization